OpenStudy (anonymous):

can someone help me write an equation of the line that is parallel to the line 2x-3y= 0 and contains the point (6,1).

OpenStudy (anonymous):

the required line will have a form exactly the line it parallel to. it means it will form 2x -3y =d (not 0) , then you plug (x =6, y =1) into new equation to find out d at the end up, put that d into the form 2x-3y =d . you get what you want